OpenStudy (anonymous):

please help!! Find the slope of a line parallel to x-y=16. 1 -1 1/16 16

OpenStudy (anonymous):

A parallel line is going to have the same slope so what is the slope of x - y = 16?

OpenStudy (anonymous):

Turn x - y = 16 into slope intercept form and that will give you your slope you need to find

OpenStudy (anonymous):

Slope intercept form look like this y = mx + b where m = slope

OpenStudy (anonymous):

i don't know but since you've already explained it can you please show me

OpenStudy (anonymous):

To change x - y = 16 into slope intercept form x - y = 16 -y = -x + 16 -y/-1 = -x/-1 + 16/-1 y = 1x -16 what is the slope?

OpenStudy (anonymous):

in slope intercept form y = mx + b The m = slope.

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Correct and that is the correct answer

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

YW